OpenStudy (anonymous):

Will fan and medal Simplify the expression 2u^2v^2 -------- 22u I know it's 1 ----- 11

OpenStudy (photon336):

\[\frac{ 2u^{2}2v^{2} }{ 22u }\]

OpenStudy (photon336):

let me ask you @WhatEven based on the problem we did before, how would you go about answering this?

OpenStudy (anonymous):

Before I would have taken it so the top read \[1uv^2\]

OpenStudy (anonymous):

(There is no 2 infront of the v)

OpenStudy (photon336):

great

OpenStudy (photon336):

what about the bottom

OpenStudy (anonymous):

But I know that that's incorrect so now I feel like it would stay the same.

OpenStudy (photon336):

because both terms are multiplied by 2 in the numerator we can factor out a 2 right? take a look here. \[\frac{ 2u^{2}2v^{2} }{ 22u } = \frac{ 2(u^{2}v^{2}) }{ 22u }\]

OpenStudy (photon336):

you already know that we can further simplify this to \[\frac{ u^{2}v^{2} }{ 11u }\] but we're not done yet. we can simplify this yet again can you see why?

OpenStudy (photon336):

@WhatEven

OpenStudy (anonymous):

Uhmmm \[\frac{ uv^2 }{11 }\]

OpenStudy (photon336):

what's your justification?

OpenStudy (anonymous):

you subtract the bottom u from the top one

OpenStudy (photon336):

exactly